THEORETICAL STUDIES OF THE ROLE OF VIBRATIONAL EXCITATION ON THE DYNAMICS OF THE HYDROGEN-TRANSFER REACTION OF F(^{2}P) + HCl → FH + Cl(^{2}P)

\underline{SARA E. RAY}, GÉ W. M. VISSERS and ANNE B. McCOY, Department of Chemistry, The Ohio State University, Columbus, OH 43210.

Hydrogen-transfer reactions are probed through vibrational excitation of the HCl bond in the pre-reactive F⋯HCl complex. Such open-shell species provide a challenge for quantum dynamical calculations due to the need to take into account multiple potential energy surfaces to accurately describe the system. A three-dimensional, fully-coupled potential energy surface has been constructed based on electronic energies calculated at the multireference configuration interaction+Davidson correction (MRCI+Q) level of theory with an aug-cc-pVnZ (n = 2, 3, 4) basis.^{a} Spin orbit calculations have also been included.^{b} Here we present the results of time-dependent quantum wave packet calculations on the asymmetric hydrogen-transfer reaction of F(^{2}P) + HCl. In these calculations, the reaction is initiated by vibrationally exciting the HCl stretching motion in the pre-reactive F⋯HCl complex. The wave packet is propagated on the coupled potential energy surfaces. Product state distributions were calculated for reactions initiated in the first three vibrationally excited states of HCl, v = 1 − 3.

aM. P. Deskevich, M. Y. Hayes, K. Takahashi, R. T. Skodje, and D. J. Nesbitt J. Chem. Phys. \underline{124}(22) 224303 (2006)

bM. P. Deskevich and D. J. Nesbitt private communication(2007)